OpenStudy (anonymous):

SOLVE THE QUAD. EQUATON BY USING THE SQUARE ROOT PROPERTY 2X^2=32

jimthompson5910 (jim_thompson5910):

\[\large 2x^2=32\] \[\large x^2=\frac{32}{2}\] \[\large x^2=16\] \[\large x=\pm\sqrt{16}\] \[\large x=\pm4\] \[\large \text{So the solutions are } x=4 \ \text{ or } \ x = -4\]

OpenStudy (anonymous):

x^2=16 then x=+4 or x=-4